

********************************************************************
UARS WEEKLY STATUS REPORT - 04 DEC 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
PEM/ZEPS was active between October 16 and October 30; PEM/AXIS
was active from November 16 to November 19 and November 27 to December 1.
During these times, PEM/ZEPS computed energy deposition from in situ electrons
and protons; PEM/AXIS computed energy deposition from an inversion of
measured X-rays.
PEM is currently conducting tests of the OBC activation. PEM activation
occurs during daylight only, when there is adequate power generated from
the UARS solar array. Currently, the PEM magnetometer and medium energy
spectrometers (particles below 30 keV) activate correctly under OBC control.
Tests conducted on the high energy particle spectrometers (HEPS -- particles
greater than 30 keV) have had limited success, and are still failing to enable
the instrument correctly. Instrument high voltage correctly activates as a
stand-alone function, but low voltage microprocessor adjustments generate
instrument errors. Additional tests are scheduled. HEPS science data between
October 1 and November 15 should be considered invalid and not used.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
After MLS was reactived successfully on Nov 21 a spacecraft power shortage
forced MLS to reduce power to allow the batteries to reach full state of
charge. On Nov 22 network problems at Goddard prevented monitoring of the MLS
instrument. After these problems persisted for several days the MLS Operations
team was forced to power off MLS early to avoid any unnecessary risk to the
instrument.
The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days HAL (4.80) PROD (L2-L3) (89R-62C=27L)
0 days HAL (4.80) PROD (L3) (1874R-1429C=445L)
41 days SOL (4.62) PROD (L2-L3) (2313R-2152C=161L)
270 days SUS (7.00) PROD (L1) (2540R-777C-1F=1762L)
6 days CLA (7.00) ENG (L1-L3) (24R-21C=3L)
UCSS developers are in the process of loading database archival information for
SOLSTICE to determine any correlation factors in SOLSTICE processing time
degradation.
CDHF preventive maintenance for the systems and peripherals is due in December.
It is scheduled for Tuesday 12/15/98.
UARS CDHF completed Y2K Time Shift testing on a subset of the
UARS CDHF Cluster. UCSS RSS services succesfully passed Y2k testing.
The CDHF is in the configuration phase of the UOAS, UCSS, MSS, and NRT
installation. This phase should not require machine downtime.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 21 November 1998 (Orbit 39299, GMT Day 325) through 26
November 1998 (Orbit 39417, GMT Day 331).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 325/0124 39314 On (PMON)
MLS 325/0124 39314 On
HRDI 325/1916 39325 Off
WINDII 325/2049 39326 Off
HRDI 327/1735 39354 On (PMON)
HRDI 328/0959 39364 Off
MLS 329/0324 39375 Off
WINDII 329/1528 39383 On
PEM 331/0006 39403 AXIS On
The beta angle ranged between 73.1 degrees to 65.8 degrees for this report
period. The beta angle is now decreasing toward a minimum angle of -1.0
degrees on 13 December 1998 (DOY 347).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.2 degrees C. Additional battery performance data
for each battery on 20 November 1998 and 27 November 1998 are provided in
Appendix D.
The clock error ranged from -10.2 to 12.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 02
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 06
% MLS 07
% PEM 02
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 08
Nine (9) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496917
Transponder B: 98/300 - 2287.498980
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/330 show the observatory data loss to be
21 hours, 39 minutes, 40.020 seconds (an increase of 24 minutes and 50.752
seconds since last report period). This is a 0.0344 percent data loss
which equals a 99.9656 percent data capture for the mission.
The increase of 24 minutes and 50.752 seconds occurred over six (6) days:
7 1.632 seconds on Day 324 due to hits in data (IP);
7 0.800 seconds on Day 325 due to hits in data (IP);
7 24 minutes and 37.120 seconds on Day 326 due to IP dropout and
insufficient overlap;
7 1.248 seconds on Day 327 due to hits in data (IP);
7 8.416 seconds on Day 328 due to hits in data (IP); and
7 1.536 seconds on Day 329 due to hits in data (IP).
The 24 minutes and 37.120 seconds data loss on Day 326 occurred in Orbit
20396. This was an added pass as the result of another IP problem in the
previous contact. Because this was an added pass, the backup stored
commands placing the recorder in record were not present. The IP problem
occurred during playback and precluded immediate corrective action by the
operators to save the downlinked data.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
WINDII has requested support for Polar Mesospheric Cloud (PMC) observations
during their next southward viewing period. WINDII is currently schedule
for observations 19-23 December 1998 and 6-10 January 1999. Both of these
periods can potentially provide PMC observations. The support required is
a roll down of the cold side by 0.5 degrees in order to lower the bottom of
WINDII's fields of view to an altitude where PMCs can be seen in the
southern hemisphere; a procedure that has been followed several times in
the past. This request is being coordinated with other UARS instruments.
The UARS FOT has started preparing the 0.5 degrees (colds side down) roll
procedure to be executed in December 1998 and January
1999.
The UARS FOT has started preparing the clock change procedure for the leap
second adjustment to be made 1 January 1999.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - No change.
Flight Dynamics Facility (FDF) - see CMS paragraph.
IP NETWORK TESTING - Complete
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The UARS FOT plans to begin testing of the Orbit Adjust Maneuver on the
UTTS(simulator) in early December 1998. Two versions will be tested; one
with B/D thrusters and another A/C thrusters. The next orbit adjust will
not be needed until late 2000 or later if the altitude limits are changed
again.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Testing of the Trending System v.1.1 release (CM-0009) continued. More
DR's and ER's are being generated.
MEETINGS
The November battery telecon was held on 24 November 1998 at 1:00 PM EST.
Only TOPEX and UARS participated. The December 1998 meeting date will not
be selected until after the Thanksgiving Holidays. The UARS batteries are
operating normally for the age of the batteries and continuing to support
the mission to 22 percent DOD with good end of night voltage.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 13
December 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 November 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39314 325/0124 98-458 PEM Power ON (PMON CNTRL)
39314 325/0124 98-459 MLS Power ON
39324 325/1728 98-460 MLS Filter Banks 2 & 4 OFF
39325 325/1916 98-461 HRDI Power OFF S/C Low Pwr
39326 325/2049 98-462 WINDII Power OFF S/C Low Pwr
39329 326/0153 98-463 MLS MLS stop scan
39333 326/0805 98-464 HRDI Reposition telescope
39333 326/0805 98-465 MLS FB 2 & 4 on, Run mode
39343 327/0014 98-466 OBC Clk Adj fr 0.00 to 20.6 msec/day
39354 327/1735 98-467 HRDI HRDI Day-Only Ops
39364 328/0959 98-468 OBC Clk Adj fr 20.6 to 41.2 msec/day
39364 328/0959 98-469 HRDI Safehold - high TBAT temp
39364 328/0959 98-470 MLS Stop scan - high SCANACT temp
39377 329/0635 98-469 HRDI Safehold - high TBAT temp
39364 329/0959 98-470 MLS Scanner OFF
39375 329/0324 98-471 MLS Power OFF
39383 329/1528 98-472 WINDII Power ON
39392 330/0641 98-469 HRDI Safehold - high TBAT temp
39403 331/0006 98-473 PEM Power AXIS ON
39408 331/0825 98-474 WINDII Error clear
39416 331/2032 98-475 OBC Clk Adj fr 40.2 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 November 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO Telescope Barrel Temp 39313-14,332-336,
361-378,391-393,
414-417
HRIFESURVTMP Y-HI Iterferometer Survival Temp 39327,328,340
HRRADTEMP Y-HI Interferometer Radiator Temp 39328
QTZEPSIF1 Y-LO I/F Plate Aft 39400
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WISCLSR WINDII Laser operating(nominal) 39395
WIERRLIMIT WINDII Software Error Code = E03 39408
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 November 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39343 327/00:27 (23 Nov 98) from 41.2 to 20.6 msec/day
39364 328/10:10 (24 Nov 98) from 20.6 to 41.2 msec/day
39416 331/20:39 (27 Nov 98) from 41.2 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 November 1998
20 Nov 1998 - GMT Day 324 Beta = 73.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +100.8/+78.4 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +16.8/-22.4 +3.87/+1.97 +18.8/-15.6 27.0 1.034 16.0
3 +78.4/+11.2 +2.60/+1.34 +18.4/-15.2 26.9 1.033 15.5
27 Nov 1998 - GMT Day 331 Beta = 65.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +95.2/+67.2 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +11.2/-33.6 +3.87/+2.60 +22.8/-15.2 26.7 1.030 17.3
3 +44.8/ 0.0 +2.92/+1.66 +22.0/-14.4 26.7 1.031 16.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 November 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-109 FOT 39323 325/1546 N/A TAC/OBC Dump problems
98-110 FOT 39325 325/1916 N/A TAC/OBC Dump problems
98-111 FOT 39341 326/2052 20396 UARS IP Desktop down
98-112 FOT 39341 326/2204 20396 IP Desktop down
98-113 FOT 39342 326/2243 N/A Commanding/AP Problems
98-114 FOT 39354 327/1735 N/A TAC/OBC Dump Problems
98-092 FOT 39407 331/0642 14170 Generic Late Acquisition
98-092 FOT 39415 331/1858 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
(NONE THIS REPORT PERIOD)
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT Nov 98 See List 14170 GENERIC LATE ACQ
98-107 FOT 39308 324/16:40 TAC 2 PROB
98-108 FOT 39309 324/17:23 TAC PROBLEMS
98-109 FOT 39323 325/15:46 TAC PROBLEMS
98-110 FOT 39325 325/19:16 TAC PROBLEMS
98-111 FOT 39341 326/20:52 20396 DESKTOP DOWN
98-112 FOT 39342 326/22:04 20396 DESKTOP DOWN
98-113 FOT 39342 326/22:43 CMD PROBLEMS
98-114 FOT 39355 327.17:35 TAC/DMP PROBLEMS
SEVNTFW48